which gas has greatest volume at stp: 0.92 moles of He 26 g of N2 7.6x 10^23 molecules of Cl2a.

CHEM 1
Answer #1

If we keep P, and T constant in ideal gas law equation then,

PV = nRT

V \alpha n

Thus the volume is directly proportional to the number of moles of gas.

To compare the volumes of given quantities of gas, we can find a number of moles and then compare.

For He, n = 0.92 moles

For N2: n = 26g/28.02g/mol = 0.9279 moles

For Cl2: n = (7.6*10^23) / (6.022*10^23) = 1.262 moles

Thus the number of moles of Cl2 is greatest and hence the volume of Cl2 will be greatest too.

The Answer is option 3) 7.6*10^23 molecules of Cl2.